ID |
Date |
Author |
Type |
Category |
Subject |
13868
|
Fri May 18 20:03:14 2018 |
Pooja | Update | Cameras | Telescopic lens solution for GigE |
Aim: To find telescopic lens solution to image test mass onto the sensor of GigE camera.
I wrote a python code to find an appropriate combination of lenses to focus the optic onto the camera keeping in mind practical constraints like distance of GigE camera from the optic ~ 1m and distance between the lenses need to be in accordance with the Thorlab lens tubes available. We have to image both the enire optic of size 3" and beam spot of 1" using this combination of lens. The image size that efficiently utilizes the entire sensor array is 1/4". Therefore the magnification required for imaging the entire optic is 1/12 and that for the beam spot is 1/4.
I checked the website of Thorlabs to get the available focal lengths of 2" lenses (instead of 1" lenses to collect sufficient power). I have tried several combination of lenses and the ones I found close enough to what is required have been listed below along with thier colorbar plots.
a) 150mm-150mm (Attachment 2 & 3)
With this combination, object distance varies like 50cm for 1" beam spot to 105cm for 3" spot. Therefore, it posses a difficulty that there is a difference of ~48cm in the distances between the optic and camera in the two cases: imaging the entire optic and the beam spot.
b) 125mm-150mm (Attachment 4 & 5)
With this combination, object distance varies like 45cm for 1" beam spot to 95cm for 3" spot. There is a difference of ~43cm in the distances between the optic and camera in the two cases: imaging the entire optic and the beam spot.
c) 125mm-125mm (Attachment 6 & 7)
The object distance varies like 45cm for 1" beam spot to 90cm for 3" spot. There is a difference of ~39cm in the distances between the optic and camera in the two cases: imaging the entire optic and the beam spot.
Sensitivity check was also done for these combination of lenses. An error of 1cm in object distance and 5mm in the distance between the lenses gives an error in magnification <2%.
The schematic of the telescopic lens system has been given in Attachment 8.
|
Attachment 1:
|
Attachment 2: plot_2018-05-18_tel-lens_150_150_1.pdf
|
|
Attachment 3: plot_2018-05-18_tel-lens_150_150_3.pdf
|
|
Attachment 4: plot_2018-05-18_tel-lens_125_150_1.pdf
|
|
Attachment 5: plot_2018-05-18_tel-lens_125_150_3.pdf
|
|
Attachment 6: plot_2018-05-18_tel-lens_125_125_1.pdf
|
|
Attachment 7: plot_2018-05-18_tel-lens_125_125_3.pdf
|
|
Attachment 8: tel_design.pdf
|
|
13867
|
Fri May 18 19:59:55 2018 |
Jon Richardson | Configuration | Electronics | AUX-PSL PLL Characterization Measurements |
Below is analysis of measurements I had taken of the AUX-PSL PLL using an SR560 as the servo controller (1 Hz single-pole low-pass, gain varied 100-500). The resulting transfer function is in good agreement with that found by Gautam and Koji (#13848). The optimal gain is found to be 200, which places the UGF at 15 kHz with a 45 deg phase margin.
For now I have reverted the PLL to use the SR560 instead of the LB1005. The issue with the LB1005 is that the TTL input for remote control only "freezes" the integrator, but does not actually reset it. This is fine if the lock is disabled in a controlled way (i.e., via the medm interface). However, if the lock is lost uncontrollably, the integrator is stuck in a garbage state that prevents re-locking. The only way to reset this integrator is to manually flip a switch on the controller box (no remote reset). Rana suggests we might be able to find a workaround using a remote-controlled relay before the controller.


|
Attachment 1: SR560_OL.pdf
|
|
Attachment 2: SR560_CL.pdf
|
|
13866
|
Fri May 18 19:10:48 2018 |
keerthana | Update | General | Code for adjusting the oscillator frequency remotly |
Target: Phase locking can be acheived by giving a scan to the oscilator frequency. This frequency is now controlled using the knobe on the AM/FM signal generator 2023B. But we need to control it remotely by giving the inputs of start frequency, end frequency and the steps.
The frequency oscilator and the computer is connected with the help of GPIB Ethernet converter. The IP address of the converter I used is '192.168.113.109' and its GPIB address is 10.
I could change the oscilator frequency by changing the input frequency with the help of the code I made (Inorder to check this code, I have changed the oscilator frequency multiple times. I hope it didn't create trouble to anyone). Now I am trying to make this code better by adding certain features like numpy, argument parse etc, which I will be able complete by next week. I am also considering to develop the code to have a sliding system to control the oscillatory frequency.
For record: The maximum limit of frequency which i changed upto is 100MHz.
|
Attachment 1: frequency_set.jpg
|
|
13865
|
Fri May 18 18:14:18 2018 |
Udit Khandelwal | Summary | General | Summary 05/18/2018 |
Tip-Tilt Suspension Design:
Designed a new ECD plate and changed dimensions of the side arms after discussing with Koji. After getting feedback on the changes, I will finish the assembly and send it to him to get approved for manufacturing.
|
13864
|
Fri May 18 14:33:34 2018 |
Kira | Update | PEM | test setup with seismometer |
Here is the result of my test. I think I'll leave the can on over the weekend because there's a long period of time where the seismometer heated up by 0.8 degrees so I can't fully see the fluctuations over a full 24 hour period. |
Attachment 1: seis_temp_can_on.png
|
|
13863
|
Fri May 18 14:18:03 2018 |
gautam | Configuration | Electronics | Basic MEDM Control Screen setup |
I setup a basic MEDM screen for remote control of the PLL.
The Slow control voltage slider allows the frequency of the laser to be moved around via the front panel slow control BNC.
The TTL signal slider provides 0/5V to allow triggering of the servo. Eventually this functionality will be transferred to the buttons (which do not work for now).
The screen can be accessed from the PSL dropdown menu in sitemap. We can make this better eventually, but this should suffice for initial setup. |
Attachment 1: AUX_PLL_CTRL.png
|
|
13862
|
Fri May 18 09:13:41 2018 |
Pooja | Update | SUS | Colored GigE image |
To obtain a colored version with good contrast of the grayscale image of scattering of light by dust particles on the surface of test mass, got using GigE camera. The original and colored images are attached here.
|
Attachment 1: Image__2017-11-14__08-25-13_100k100g1V_colored.png
|
|
Attachment 2: Image__2017-11-14__08-25-13_100k100g1V.tiff
|
13861
|
Fri May 18 07:41:01 2018 |
steve | Update | SUS | clipping ITMX oplev |
The ITMX oplev still clipping
Quote: |
The ITMX oplev beam is clipping. It will be corected with locked arm
|
|
13860
|
Thu May 17 18:05:01 2018 |
gautam | Update | SUS | SR785 near 1X5 |
I'm working near 1X5 and there is an SR785 adjacent to the electronics rack with some cabling running along the floor. I plan to continue in the evening so please leave the setup as is.
During the course of this work, I noticed the +15V Sorensen in 1X6 has 6.8 A of current draw, while Steve's February2018 label says the current draw is 8.6A. Is this just a typo?
Steve: It was most likely my mistake. Tag is corrected to 6.8A
I'm still in the process of electronics characterization, so the SR785 is still hooked up. MC3 coil driver signal is broken out to measure the output voltage going to the coil (via Gainx100 SR560 Preamp), but MC is locked. |
Attachment 1: B55CE985-B703-4282-B716-3144957C7372.jpeg
|
|
13859
|
Thu May 17 15:38:19 2018 |
Kira | Update | PEM | test setup with seismometer |
I've moved my setup to the actual seismometer. I attached the temperature sensor to the seismometer (attachment 1) with duct tape, though this is temporary. I will be monitoring the temperature fluctuations of the seismometer for a whole day then take the can off and repeat the test. The can isn't clamped down so the insulation isn't perfect, so I'd expect to see some noticeable fluctuations even with the can on. I've also labeled the long cable for the temperatuse sensor readout (attachments 2 and 3). There will also be an out of loop sensor added in later, but for this test since I am not running the loop it doesn't matter which sensor I monitor. Attachment 4 is a picture of the current setup. |
Attachment 1: IMG_20180517_144420.jpg
|
|
Attachment 2: IMG_20180517_145754.jpg
|
|
Attachment 3: IMG_20180517_151956.jpg
|
|
Attachment 4: IMG_20180517_145121.jpg
|
|
13858
|
Thu May 17 13:51:35 2018 |
Jon Richardson | Configuration | Electronics | Documentation & Schematics for AUX-PSL PLL |
[Jon, Gautam]
Attached is supporting documentation for the AUX-PSL PLL electronics installed in the lower PSL shelf, as referenced in #13845.
Some initial loop measurements by Gautam and Koji (#13848) compare the performance of the LB1005 vs. an SR560 as the controller, and find the LB1005 to be advantageous (a higher UGF and phase margin). I have some additional measurements which I'll post separately.
Loop Design
Pickoffs of the AUX and PSL beams are routed onto a broadband-sensitive New Focus 1811 PD. The AUX laser temperature is tuned to place the optical beat note of the two fields near 50 MHz. The RF beat note is sensed by the AC-coupled PD channel, amplified, and mixed-down with a 50 MHz RF source to obtain a DC error signal. The down-converted term is isolated via a 1.9-MHz low-pass filter in parallel with a 50 Ohm resistor and fed into a Newport LB1005 proportional-integral (PI) servo controller. Controller settings are documented in the below schematic. The resulting control signal is fed back into the fast PZT actuator input of the AUX laser.
Schematic diagram of the PLL.
Hardware Photos
Optical layout on the PSL table.
PLL electronics installed in the lower PSL shelf.
Close-up view of the phase detector electronics.
Slow temp. (left) and fast PZT signals into the AUX controller.
AUX-PSL beat note locked at 50 MHz offset, from the control room.
|
13852
|
Thu May 17 11:56:37 2018 |
gautam | Update | General | EPICS process died on c1ioo |
The EPICS process on the c1ioo front end had died mysteriously. As a result, MC autolocker wasn't working, since the autolocker control variables are EPICS channels defined in the c1ioo model. I restarted the model, and now MCautolocker works. |
13851
|
Thu May 17 09:14:38 2018 |
Steve | Update | General | Stack measurement setup decommissioned |
The final set-up of stack measurment with 3 load cells and 4 leveling wedge mounts as Atm 1
Sensor voltages BEFORE and AFTER this attempt. |
Attachment 1: Load_Cell_Measurement_Set_Up.jpg
|
|
Attachment 2: ETMX_stack_up_down.png
|
|
13850
|
Wed May 16 21:47:17 2018 |
Kevin | Update | PEM | Seismometer Noise Spectra |
Earlier today Kira and I reconnected the EX seismometer. I just took some spectra of all three seismometers, shown in the attachments, to compare with past data and to do a rough check of the calibration.
This elog has a spectra from 2010 (GUR1 is now EY) and this elog has one for BS at lower frequencies from 2017. Note that the EX seismometers now have strong peaks that are not at 60 Hz harmonics. Other than these peaks, these old spectra roughly match up with the ones taken today, so the callibration is still roughly the same. I couldn't find any old data for EX (GUR2) though so I don't know for sure that these peaks weren't there before.
gautam 20180517 0930: In 2017, Gur2 (now EX) looked like this. Still peaky, but the peaks seem shifted in frequency. Steve also informed me that the Gur1 and Gur2 cables were swapped n times, so perhaps we shouldn't read too much into that. |
Attachment 1: BS_vel.pdf
|
|
Attachment 2: EX_vel.pdf
|
|
Attachment 3: EY_vel.pdf
|
|
13849
|
Wed May 16 21:02:22 2018 |
Kevin | Update | PEM | ADC common mode rejection with new seismometer connections |
As described in this elog, the ADC for the seismometers now has the signals wired directly to the ADC instead of going through an AA board or other circuit to remove any common mode noise. This elog describes one test of the common mode rejection of this setup. Guantanamo suggested comparing directly with a recent spectrum taken a few months before the new setup described in this elog.
Today I took a spectrum (attachment 1) of C1:PEM-MIC_2 (Ch17) and C1:PEM-MIC_3 (Ch18) with input to the ADC terminated with 50 Ohms. These are two of the channels plotted in the previous spectrum, though I don't know how that plot was normalized. It's clear that there are now strong 60 Hz harmonic peaks that were not there before, so this new setup does have worse common mode rejection. |
Attachment 1: ADC_noise.pdf
|
|
13848
|
Wed May 16 18:52:50 2018 |
gautam | Configuration | Electronics | PLL mysteries solved |
[Koji, Gautam]
Summary:
As I suspected, when the SR560 is operated in 1 Hz, first order LPF mode, the (electronic) transfer function has a zero at ~5kHz (!!!).
Details:
This is what allowed the PLL to be locked with this setting with UGF of ~30kHz. On the evidence of Attachment #3, there is also some flattening of the electrical TF at low frequencies when the SR560 is driving the NPRO PZT. I'm pretty sure the flattening is not a data download error but since this issue needs further investigation anyway, I'm not reading too much into it. I fit the model with LISO but since we don't have low frequency (~1Hz) data, the fit isn't great, so I'm excluding it from the plots.
We also did some PLL loop characterization. We decided that the higher output range (10Vp bs 10Vpp for the SR560) of the LB1005 controller means it is a better option for the PLL. The lock state can also be triggered remotely. It was locked with UGF ~ 60kHz, PM ~45deg.
We also measured the actuation coefficient of the NPRO laser PZT to be 4.89 +/- 0.02 MHz/V. Quoted error is (1-sigma) from the fit of the linear part of the measured transfer function to a single pole at DC with unknown gain. I used the "clean" part of the measurement that extends to lower frequencies for the fit, as can be seen from the residuals plot. Good to know that even though the LDs are dying, the PZT is still going strong :D.
Remaining loop characterization (i.e. verification of correct scaling of in loop suppression with loop gain etc.) is left to Jon.
Measurement schemes:
- OLG (Attachment #1) was measured using the usual IN1/IN2 technique.
- PZT calibration (Attachment #2) was measured by injecting an excitation at the PLL control point.
- The ratio of the PLL error point (Volts) to Excitation (Volts) was measured using the SR785.
- The error point was calibrated by looking at the PLL open loop Vpp (corresponds to pi radians of phase shift).
- Dividing the fitted gain of the phase->Frequency conversion by the error point calibration, we get the PZT actuation coefficient.
Some other remarks:
- In the swept-sine mode, the SR785 measures transfer functions by taking the ratio of complex FFT values of its inputs at the drive frequency. So the phase in particular is a good indicator of whether the measurement is coherent or not.
- In all these measurements, the PLL gain is huge at low frequencies, and hence, the excitation is completely squished on propagating through the loop. E.g. a 10mV excitation is suppressed by a factor of ~60dB = 1000 to 10uV, and if the analyzer autoRange is set to UpOnly, it is easy to see how this is drowned at the IN1 input. This is why the measurements lose coherence below ~1 kHz.
- It is easy to imagine implementing an EPICS servo that offloads the DC part of the LB box control signal to the SLOW frequency input on the Lightwave controller. This would presumably allow us to extend the lock timescales. We can also easily implement a PLL autolocker.
- Preliminary investigation of the SR560 situation suggests that individual filter stages can only achieve a maximum stopband attenuation of 60dB relative to the passband. When we cascade two stages together, 120dB seems possible...
|
Attachment 1: PLLanalysis.pdf
|
|
Attachment 2: PZTcal.pdf
|
|
Attachment 3: SR560_funkiness.pdf
|
|
13847
|
Tue May 15 22:11:38 2018 |
gautam | Update | General | IFO maintenance |
Since there have been various software/hardware activity going on (stack weighing, AUX laser PLL, computing timing errors etc etc), I decided to do a check on the state of the IFO.
- c1susaux, c1aux and c1iscaux crates were keyed as they were un-telnet-able.
- Single arm locking worked fine, TT alignment was tweaked (as these had drifted due to the ADC failure in c1lsc) to maximize Y arm transmission using the dither servos.
- Arms weren't staying locked for extended periods of time. I particularly suspected ITMX, as I saw what I judged to be excess motion on the Oplev.
- @Steve - ITMX and BS HeNes look like they are in need of replacement judging by the RIN (although the trend data doesn't show any precipitous drop in power). If we are replacing the BS/PRM Oplev HeNe, might be a good time to plan the inejction path a bit better on that table.
- RIN in Attachment #1 has been normalized by the mean value of the OL sum channel. There is now a script to make this kind of plot from NDS in the scripts directory (as I found it confusing to apply different calibrations to individual traces in DTT).
|
Attachment 1: OL_RIN_2018_05_15.pdf
|
|
Attachment 2: OLsums.png
|
|
13846
|
Tue May 15 21:56:57 2018 |
gautam | Update | General | Stack measurement setup decommissioned |
[steve,koji,gautam]
Since we think we already know the stack mass to ~25% (i.e. 5000 +/- 1000 lbs), we decided to restore the ETMX stack. Procedure followed was:
- Take photos of all dial indicators and spirit level. We were at ~-22 mils on all 3 indicators, with 0 being the level before we touched the stack two Fridays ago, i.e. May4.
- Raised all four jacks installed underneat blue crossbeams in 5mil increments until we were at +25mils on all of them. At this point, there was negligible load on the load cells on top of the STACIS legs, and we could easily slide the load cells out.
- Rotated all jack screws clockwise (i.e. moving jack screws downwards) by 270 degrees. The southeast jackscrew was rotated by an additional 360 degrees. This was to undo all the jack-screw raising we did on Friday, May 4.
- Re-installed jacks which were present originally on the STACIS legs, taking care to center the jack as best as we could by eye on the STACIS leg, per Dennis Coyne's suggestion not to impose shear strain on STACIS legs. There were supposedly never carrying any load, and are according to Steve, are there more for safety purposes.
- Lowered all four jacks in 5 mil steps until dial indicators read ~0. The Northwest jack resting on the STACIS leg was somehow ~0.5cm (!!) below the blue crossbeam even though the corresponding dial gauge read 0, so we raised the jack until it was barely grazing the bottom of the blue crossbeam (confirmed by looking at the point where the dial indicator started going up again). Not sure why this should have been, best hypothesis we have is that someone (one of us) changed the level of this jack while it was removed from the setup.
- Checked that jack screws could not be turned by hand. At this point, all the load has to be resting on the jack screws, as the jacks we had installed to raise the blue crossbeams could be slid out from underneath the blue beams and hence were carrying no load.
- Took photographs of all dial indicators, spirit level. We were satisfied that we had recovered the "nominal" stack alignment as best as we could judge with the available indicators.
- ETMX Oplev spot had returned to the PD. ETMX watchdog was re-engaged, optic was re-aligned using SLOW bias sliders to center Oplev spot.
- EX NPRO was turned back on, and the green beam was readily locked to a cavity TEM00 mode
.
I will upload the photos to the PICASA page and post the link here later.
Quote: |
In this case, we only need a mass estimate of the end chamber contents with an accuracy of ~25%. If we think we have that already, we don't need to keep doing the jacks-strain gauge adventure.
|
|
13845
|
Tue May 15 20:51:27 2018 |
gautam | Configuration | Electronics | Making PLL setup more permanent |
[jon, steve, gautam]
Some points which Jon will elaborate upon (and put photos of) in his detailed elog about this setup:
- PLL electronics (mixer, coupler, ZFL500HLN amplifier and DC power supply for the beatnote, SR560 servo) all reside on the newly installed lower level PSL shelf.
- Cross connect channel C1:PSL-126MOPA_126CURADJ hijacked for remote temperature control of the AUX NPRO. Note that shield of front panel BNC is ground and so even though the manual says the controller accepts +/-10V, this is not a differential input. BNC cable was routed from cross connect to PSL enclosure, MEDM slider will be setup.
- There was an SMA cable running from the VEA to the control room which we decided to use for monitoring of the beatnote amplitude on the control room analyzer. Yesterday, Steve and I routed the end of this inside the VEA, near 1X2 originally, to inside the PSL table where it is hooked up to the (20dB) coupled amplifier output. This required some work on the cable tray, we were careful but in case there is some wonkiness in some signals, perhaps this work is to blame.
We are now in a state where the PLL can be locked remotely from the control room by tweaking the AUX laser temperature . Tomorrow, Keerthana will work on getting Craig's/Johannes' Digital Frequency Counter script working here, I think we can easily implement a PLL autolocker if we have some diagostic that tells us if the PLL us locked or not.
Steve informed me that there is an acoustic hum inside the PSL enclosure which wasn't there before. Indeed, it is at ~295Hz, and is from the Bench power supply used to power the ZFL500HLN amplifier. This will have to go... |
13844
|
Tue May 15 15:13:23 2018 |
Kira | Summary | PEM | Acromag issues |
I tried calibrating the other channels today, but they still fluctuate. Sometimes they do stabilize at +/- 10V, but then suddenly drop to 5 or 6 V before climbing back up to 10. Turning the legacy off made it go only up to 6.67V. This happens for all the channels, even after doing a factory reset and recalibrating. Not sure what's happening here. |
13843
|
Tue May 15 13:34:30 2018 |
Steve | Update | safety | surf safety training |
Pooja and Keirthana received 40m specific basic safety training. |
13842
|
Tue May 15 10:42:14 2018 |
Koji | Update | PSL | Modulation depth measurement for the 3IFO aLIGO EOM and aftermath |
The marconi RF output was turned on and thus the RF generator condition was restored to the nominal state on Friday 11th. |
13841
|
Mon May 14 18:58:32 2018 |
Kevin | Update | PonderSqueeze | Squeezing with no SRM |
Quote: |
Note that for Signal Recycling, which is what Kevin tells us we need to do, there is a DARM pole at ~150 Hz.
|
To be quantitative, since we are looking at smaller squeezing levels and considering the possibility of using 5 W input power, it is possible to see a small amount of squeezing below vacuum with no SRM.
Attachment 1 shows the amount of squeezing below vacuum obtainable as a function of homodyne angle with no SRM and 5 W incident on the back of PRM. The optimum homodyne angle at 210 Hz is 89.2 deg which gives -0.38 dBvac of squeezing. Figure 2 is the displacement noise at this optimal homodyne angle and attachment 3 is the same noise budget shown as the ratio of the various noise sources to the unsqueezed vacuum.
The other parameters used for these calculations are:
- 4.5 kΩ series resistance for the ETM coils; 15 kΩ for the ITM coils
- 100 ppm transmissivity on the folding mirrors giving a PRC gain of 40
- PD quantum efficiency of 0.88
So maybe it's worth considering going for less squeezing with no SRM if that makes it technically more feasible. |
Attachment 1: homodyne_heatmap.pdf
|
|
Attachment 2: displacement_noise.pdf
|
|
Attachment 3: noise_budget.pdf
|
|
13840
|
Mon May 14 08:55:40 2018 |
Dennis Coyne | HowTo | SEI | preparation of load cell measurement at ETMX |
follow up email from Dennis 5-13-2018. The last line agrees with the numbers in elog13821.
Hi Steve & Gautam,
I've made some measurements of the spare (damaged) 40m bellows. Unfortunately neither of our coordinate measurement arms are currently set up (and I couldn't find an appropriate micrometer or caliper), so I could not (yet) directly measure the thickness. However from the other dimensional measurements, and a measurement of the axial stiffness (100 lb/in), and calculations (from the Standards of the Expansion Joint Manufacturers Association (EJMA), 6th ed., 1993) I infer a thickness of 0.010 inch in . This is close to a value of 0.012 in used by MDC Vacuum for bellows of about this size.
I calculate that the maximum allowable torsional rotation is 1.3 mrad. This corresponds to a differential height, across the 32 in span between support points, of 0.041 in.
In addition using the EJMA formulas I find that one can laterally displace the bellows by 0.50 inch (assuming a simultaneous axial displacement of 0.25 inch, but no torsion), but no more than ~200 times. I might be good to stay well below this limit, say no more than ~0.25 inch (6 mm).
If interested I've uploaded my calculations as a file associated with the bellows drawing at D990577-A/v1.
BTW in some notes that I was given (by either Larry Jones or Alan Weinstein) related to the 40m Stacis units, I see a sketch from Steve dated 3/2000 faxed to TMC which indicates 1200 lbs on each of two Stacis units and 2400 on the third Stacis.
|
13839
|
Sun May 13 20:48:38 2018 |
johannes | Update | General | CDS crash |
I think the root of the problem is that the /opt/rtapps/ and /cvs/cds/rtapps/ mounting locations point to the same directory on the nfs server. Gautam and I were cleaning up the /cvs/cds/caltech/target/ directory, placing the previous contents of /cvs/cds/caltech/target/c1auxex/, including database files and startup instructions in /cvs/cds/caltech/target/c1auxex_oldVME/, and then moved /cvs/cds/caltech/target/c1auxex2/, which has the channel database and initialization files for the Acromac DAQ, to /cvs/cds/caltech/target/c1auxex/.
This also required updating the systemd entries on c1auxex to point to the changed directory. While confirming that everything worked as before we noticed that upon startup the EPICS IOC complains about not being able to find the caRepeater binary. This was not new and has not limited DAQ functionality in the past, but we wanted to fix this, as it seemed to be some simple PATH issue. While the paths are all correctly defined in the user login shell, systemd runs on a lower level and doesn't know about them. One thing we tried was to let systemd execute /cvs/cds/rtapps/epics-3.14.12.2_long/etc/epics-user-env.sh initializing EPICS. It was strange that the content of that file was pointing to /opt/rtapps/epics-3.14.12.2_long/base, which is not mounted on the slow machines, so we changed the /opt/ it to /cvs/cds/, not realizing that the frontends read from the same directory (as Gautam said, /cvs/cds does not exist as a mount point on the frontend). It ended up not working this way, and apparently I forgot to change it back during clean up. But worse, never elogged it!
In the end, we managed to to give systemd the correct path definitions by explicitly calling them out in /cvs/cds/caltech/target/c1auxex/ETMXenv, to which a reference was added in the systemd service file. The caRepeater warning no longer appears. |
13838
|
Sun May 13 17:31:51 2018 |
gautam | Update | General | CDS crash |
As suspected, this was indeed a path problem. Johannes will elog about it later, but in short, it is related to some path variables being changed in order to try and streamline the EPICS processes on the new c1auxex machine (Acromag Era). It is confusing that futzing around with the slow computing system messes with the realtime system as well - aren't these supposed to be decoupled? Once the paths were restored by Johannes, everything compiled and restarted fine. We even have a beam on the AS camera, which was what triggered this whole thing .
Anyways, Attachment #1 shows the current status. I am puzzled by the red TIMING indicators on the c1x04 and c1x02 processes, it is absent from any other processes. How can this be debugged further?
Quote: |
I suspect this is some kind of path problem - the EPICS_BASE bash variable is set to /cvs/cds/rtapps/epics-3.14.12.2_long/base on the FEs, while /cvs isn't even mounted on the FEs (nor do I think it should be). I think the correct path should be /opt/rtapps/epics-3.14.12.2_long/base. Why should this have changed?
|
|
Attachment 1: CDS_overview_20180513.png
|
|
Attachment 2: AS_1210293643.jpeg
|
|
13837
|
Sun May 13 15:15:18 2018 |
gautam | Update | General | CDS crash |
I found the c1lsc machine to be completely unresponsive today. Looking at the trend of the state word, it happened sometime yesterday (Saturday). The usual reboot procedure did not work - I am not able to bring back any of the models on any of the machines, during the restart procedure, they all fail. The logfile reads (for the c1ioo front end, but they all behave the same):
[ 309.783460] c1x03: Initializing space for daqLib buffers
[ 309.887357] CPU 2 is now offline
[ 309.887422] c1x03: Sync source = 4
[ 309.887425] c1x03: Waiting for EPICS BURT Restore = 2
[ 309.946320] c1x03: Waiting for EPICS BURT 0
[ 309.946320] c1x03: BURT Restore Complete
[ 309.946320] c1x03: Corrupted Epics data: module=0 filter=1 filterType=0 filtSections=134610112
[ 309.946320] c1x03: Filter module init failed, exiting
[ 363.229086] c1x03: Setting stop_working_threads to 1
[ 364.232148] DXH Adapter 0 : BROADCAST - dx_user_mcast_unbind - mcgroupid=0x3
[ 364.233689] Will bring back CPU 2
[ 365.236674] Booting Node 1 Processor 2 APIC 0x2
[ 365.236771] smpboot cpu 2: start_ip = 9a000
[ 309.946320] Calibrating delay loop (skipped) already calibrated this CPU
[ 365.251060] NMI watchdog enabled, takes one hw-pmu counter.
[ 365.252135] Brought the CPU back up
[ 365.252138] c1x03: Just before returning from cleanup_module for c1x03
Not sure what is going on here, or what "Corrutped EPICS data" is supposed to mean. Thinking that something was messed up the last time the model was compiled, I tried recompiling the IOP model. But I'm not able to even compile the model, it fails giving the error message
make[1]: Leaving directory '/opt/rtcds/caltech/c1/rtbuild/3.4'
make[1]: /cvs/cds/rtapps/epics-3.14.12.2_long/modules/seq/bin/linux-x86_64/snc: Command not found
make[1]: *** [build/c1x03epics/c1x03.c] Error 127
Makefile:28: recipe for target 'c1x03' failed
make: *** [c1x03] Error 1
I suspect this is some kind of path problem - the EPICS_BASE bash variable is set to /cvs/cds/rtapps/epics-3.14.12.2_long/base on the FEs, while /cvs isn't even mounted on the FEs (nor do I think it should be). I think the correct path should be /opt/rtapps/epics-3.14.12.2_long/base. Why should this have changed?
I've shutdown all watchdogs until this is resolved. |
Attachment 1: vertexFEs_crashed.png
|
|
13836
|
Sat May 12 10:02:03 2018 |
rana | Update | General | More refinement of DARM control signal projection |
Good question! I've never calculated what the resonance frequency would be if had an inductive resistor with our cable capacitance (~50 pF/m I guess). |
13835
|
Fri May 11 19:02:52 2018 |
gautam | Update | General | More refinement of DARM control signal projection |
I was a bit hasty in posting the earlier plots. In the earlier plot, the "OLG" trace was OLG * anti dewhitening as Rana pointed out.
Here are the updated ones, and a cartoon (Attachment #5) of the loop topology I assumed. I've excluded things like violin filters, AA/AI etc. The overall gain scaling I mentioned in the previous elog amounts to changing the optical sensing response in this cartoon. I now also show the DARM suppression (Attachment #4) for this OLG and the DARM linewidths for RSE. I don't think the conclusions change.
Note that for Signal Recycling, which is what Kevin tells us we need to do, there is a DARM pole at ~150 Hz. I assume we will cancel this in the digital controller and so can achieve a similar OLG shape. This would modify the control signal spectrum a little around 150Hz. But for a UGF on the loop of ~150 Hz, we should still be able to roll-off the control signal at high frequencies and so the RMS shouldn't be dramatically affected.
Steve is looking into acquiring 4.5kohm Vishay Wirewound resistors with 1% tolerance. Plan is to install two in parallel (so that we get 2kohm effective resistance) and then snip off one once we are convinced we won't have any actuation range issues. Do these look okay? They're ~$1.50ea on mouser assuming we get 100. Do we need the non-inductive winding?
Quote: |
I think "OLG" trace is not labeled right; it would be good to see the actual OLG in addition to whatever that trace actually is.
|
|
Attachment 1: darmProj.pdf
|
|
Attachment 2: darmOLTF.pdf
|
|
Attachment 3: DWcomparison.pdf
|
|
Attachment 4: DARMsuppression.pdf
|
|
Attachment 5: ControlLoop.pdf
|
|
13834
|
Fri May 11 18:17:07 2018 |
gautam | Update | DetChar | AUX laser PLL setup |
[koji, gautam]
As discussed at the meeting earlier this week, we will use some old *MOPA* channels for interfacing with the PLL system Jon is setting up. He is going to put a sketch+photos up here shortly, but in the meantime, Koji helped me identify a channel that can be used to tune the temperature of the Lightwave NPRO crystal via front panel BNC input. It is C1:PSL-126MOPA_126CURADJ, and is configured to output between +/-10V, which is exactly what the controller can accept. The conversion factor from EPICS value to volts is currently set to 1 (i.e. EPICS value of +1 corresponds to +1V output from the DAC). With the help of the wiring diagram, we identified pins 3 and 4 on cross-connect #J7 as the differential outputs corresponding to this channel. Not sure if we need to also setup a TTL channel for servo ENABLE/DISABLE, but if so, the wiring diagram should help us identify this as well.
The cable from the DAC to the cross-connect was wrongly labelled. I fixed this now. |
13833
|
Fri May 11 13:58:42 2018 |
rana | Update | General | More refinement of DARM control signal projection |
I think "OLG" trace is not labeled right; it would be good to see the actual OLG in addition to whatever that trace actually is.
Based on the first plot, however, my conclusion is that:
- we don't need the passive isolator to reduce the control signal; the control signal is dominated by f < 10 Hz.
- we should still look into isolators for the reduction of the f > 50 Hz stuff, just to make the overall DARM sensitivity better. But this does not have to be pneumatic since we no longer need 10 Hz isolation. It can instead be a solid piece of rubber to give us a ~20-30 Hz resonance. That would still give us a factor of 5-10 improvement above 100 Hz.
- In this case, we only need a mass estimate of the end chamber contents with an accuracy of ~25%. If we think we have that already, we don't need to keep doing the jacks-strain gauge adventure.
|
13832
|
Fri May 11 11:47:33 2018 |
johannes | Summary | PEM | Acromag issues |
The replacement Acromag we scooped from the West Bridge E-Shop does actually seem to work, although we thought it was broken - at first it was just outputting zeros, but after I did the calibration procedure, applying +10 V and -10 V, respectively, it was reporting voltage correctly, over the full range. I don't know why the factory settings would be messed up, but it had been out of the box before. I did this only with channel 7, so you need to calibrate channels 0-6 and confirm that they indeed also work properly. |
13831
|
Thu May 10 14:13:22 2018 |
gautam | Update | General | More refinement of DARM control signal projection |
Summary:
- It seems that after a x10 increase in the coil driver resistance, we will have enough actuation range to control (anti de-whitened) DARM without saturating the DAC.
- The Barry puck doesn't seem to help us much in reducing the required RMS for DARM control. If this calculation is to be believed, it actually makes the RMS actuation a little bit higher.
See Attachment #1 for the projected control signal ASDs. The main assumption in the above is that all other control loops can be low-passed sufficiently such that even with anti-dewhitening, we won't run into saturation issues.
DARM control loop:
- I'm now calculating the DARM control signal in counts after factoring into account a digital DARM control loop.
- The loop shape is what we used when the DRFPMI was locked in Oct 2015.
- I scaled the overall OLTF gain to have a UGF around 200Hz.
- The breakdown of how the DARM loop is constructed is shown in Attachment #2.
De-whitening and Anti-De-whitening:
- The existing DW shape in the ITM and ETM signal chains has ~80dB attenuation around 100 Hz.
- Assuming ~5uV/rtHz noise from the DAC, 60dB of low-passing gets us to 5nV/rtHz. With 4.3kohm series resistance, this amounts to ~1pA/rtHz current noise (compared to ~3pA/rtHz from the Johnson noise of the series resistance). Actually, I measured the DAC noise to be more like ~700nV/rtHz at 100 Hz, so the current noise contribution is only 0.16pA/rtHz.
- This amounts to getting rid of the passive filter at the end of the chain in the de-whitening board.
- Attachment #3 shows the existing and proposed filter shapes.
It remains to add the control signals for Oplev, local damping, and ASC to make sure we have sufficient headroom, but given that current projections are predicting using up only ~1000cts of the ~23000cts (RMS) available from the DAC, I think it is likely we won't run into saturations. Need to also figure out what the implication of the reduced actuation range will be on handling the locking transient. |
Attachment 1: darmProj.pdf
|
|
Attachment 2: darmOLTF.pdf
|
|
Attachment 3: DWcomparison.pdf
|
|
13830
|
Thu May 10 11:38:19 2018 |
gautam | Update | General | ITMY UL |
Looking at Steve's plot, I was reminded of the ITMY UL OSEM issue. The numbers don't make sense to me though - 300um of DC shift in UL with negligible shifts in the other coils should have made a much bigger DC shift in the Oplev spot position. |
Attachment 1: ITMY_UL.pdf
|
|
13829
|
Thu May 10 08:45:16 2018 |
Steve | Update | General | 4.5M eq. Cabazon, CA |
20180508 4:49am Cabazon earth quake 4.5M at 79 miles away. ETMX is in load cell measurment condition.
Quote: |
There was an earthquake, all watchdogs were tripped, ITMX was stuck, and c1psl was dead so MCautolocker was stuck.
Watchdogs were reset (except ETMX which remains shutdown until we finish with the stack weight measurement), ITMX was unstuck using the usual jiggling technique, and the c1psl crate was keyed.
|
|
Attachment 1: Cabazon4.5m79m.png
|
|
Attachment 2: 4.5Meq.png
|
|
13828
|
Wed May 9 19:51:07 2018 |
gautam | Update | General | Input beam misaligned |
As suspected - the problem was with the TTs. I tested the TT signal chain by driving a low frequency sine wave using AWG and looking at the signal on an o'scope. But I saw nothing, neither at the AI board monitor point, nor at the actual coil driver mon point. I decided to look at the IOP testpoints for the DAC channels, to see if the signals were going through okay on the digital side. But the IOP channels were flatlined, as viewed on dataviewer (see Attachment #1). This despite the fact that the DAC output monitor screen in the model itself was showing some sensible numbers, again see Attachment #1.
Looking at the CDS overview screen, there were no red flags. But there was a red indicator sneakily hidden away in the IOP model's CDS status screen, the "DAC" field in the state word is red. As Attachment #2 shows, a change in the state word is correlated with the time ASDC went to 0.
Note that there are also no errors on the c1lsc frontend itself, judging by dmesg. I want to do a model restart, but (i) this will likely require reboots of all vertex FEs and (ii) I want to know if any CDS experts want to sniff for clues to what's going on before a model restart wipes out some secret logfiles. I'm a little confused that the rtcds isn't throwing up any errors and causing models to crash if the values are not being written to the registers of the DAC. It may also be that the DAC card itself is dead . To re-iterate, all the EPICS readbacks were suggesting that I am injecting a signal right up to the IOP.
Quoting from the runtime diagnostics note:
NOTE: As V2.7, if this error is detected, the IOP will output zero values to all DAC modules, as a protective measure. Only method to clear this is to restart the IOP and all applications on that computer
|
Attachment 1: DACweirdness.png
|
|
Attachment 2: DACerror.png
|
|
13827
|
Wed May 9 17:30:04 2018 |
gautam | Update | General | Input beam misaligned |
There is no beam going into the IFO at the moment. There was definitely a spot on the AS camera after I restored the suspensions yesterday, as you can see from the ASDC level in Attachment #1. But at around 2pm Pacific yesterday, the ASDC level has gone to 0. I suspect the TTs. There is no beam on the REFL camera either when PRM is aligned, and PRM's DC alignment is good as judged by Oplev.
Normally, I am able to recover the beam by scanning the TTs around with some low frequency sine waves, but not today. We don't have any readback (Oplev/OSEM) of the TT alignment, and the DC bias values havent jumped abnormally around the time this happened, judging by the OUT16 monitor points (see Attachment #2). The IMC was also locked at the time when this abrupt drop in ASDC level happened. Unfortunately, we don't have a camera on the Faraday so I don't know where the misalignment is happening, but the beam is certainly not making it to the BS. All the SOS optics (e.g. BS, ITMX and ITMY) are well aligned as judged by Oplev.
Being debugged now... |
Attachment 1: InputBeamGone.png
|
|
Attachment 2: TTpointing.png
|
|
13826
|
Tue May 8 11:41:16 2018 |
gautam | Update | General | IFO maintenance |
There was an earthquake, all watchdogs were tripped, ITMX was stuck, and c1psl was dead so MCautolocker was stuck.
Watchdogs were reset (except ETMX which remains shutdown until we finish with the stack weight measurement), ITMX was unstuck using the usual jiggling technique, and the c1psl crate was keyed. |
Attachment 1: ITMX_stuck.png
|
|
13825
|
Tue May 8 10:24:10 2018 |
Kira | Summary | PEM | plan for this week |
Here are a few things I will be working on:
- Design PCB boards for the heater circuit and temperature sensor circuits [by wednesday]
- Order the front panel I've designed for the seismometer block [today]
- [next week?] install the new Acromag when it comes
|
13824
|
Tue May 8 00:40:51 2018 |
gautam | Configuration | ALS | BeathMouth pulled out of PSL table |
Summary:
I did some more BeatMouth characterization. My primary objective was to do a power budget. Specifically, to measure the insertion loss of the mating sleeves and of the fiber beam splitters. All power numbers quoted in this elog are measured with the fiber power meter. Main takeaways:
- Measured insertion loss of all mating sleeves, which are ADAFCPMB2, are in agreement with the < 1dB spec. 1 dB in power is ~20%.
- But there is large variance in the above number, between different supposedly identical connectors.
- Measured insertion loss from input port to coupled ports of the fiber beamsplitters are slightly larger than spec (~3.5dB), when measured after mating the fiber beamsplitter (which has Hi1060 flex fiber) and PM980 fiber (which is what brings light to the BeatMouth).
- But measured insertion loss when mating is between Hi1060 flex fiber ends is more in line with the expected value of ~3.5dB.
- Isolation of fiber beam splitters seems to match the spec of >55dB.
Results:
- I used the fiber bringing 416uW of IR light from EY for this test.
- Insertion loss was measured by injecting the fiber light at one port and measuring the transmitted power at various other ports.
- In order to couple the fiber power meter across a single mating sleeve, I used a short (~1m) patch cable from newport (F-SY-C-1FCA). Technically, this is single mode fiber with the correct type of connector, FC/APC, but is not PM.
- See Attachment #2 for the labeling of the connectors which is how I refer to them in the table below.
- Lest there be confusion, I use the definition of insertion loss
.
Mating Sleeve # |
Insertion loss [dB] |
1 |
0.38 |
2 |
0.65 |
3 |
0.71 |
4 |
0.43 |
5 |
0.95 |
6 |
0.79 |
7 |
0.5 |
Remarks / Questions:
- Is there any way to systematically reduce the insertion loss? Like getting a better mating part?
- Question for the fiber experts: How do we deal with the unused port of the beam-splitters? Right now, they are just capped. But as you can see in Attachment #1, the caps certainly don't block all the light. What are the implications of back-scattered light into the fiber on the ALS noise? I guess the beamsplitter itself has the spec'd 55dB directivity, so do we not care about this?
|
Attachment 1: IMG_6986.JPG
|
|
Attachment 2: IMG_6987.JPG
|
|
13823
|
Mon May 7 20:01:14 2018 |
Rorpheus | Update | General | Use anti-dewhitening + show CARMA/DARMA |

example of plots illustrating DAC range / saturation |
13822
|
Mon May 7 16:23:06 2018 |
gautam | Update | General | DARM actuation estimate |
Summary:
Using the Wiener Filter estimate of the DARM disturbance we will have to cancel, I computed how the control signal would look like for a few scenarios. Our DACs are 16-bit, +/-10V (i.e. +/-32,768cts-pk, or ~23,000cts RMS). We need to consider the shape of the de-whitening filter to conclude whether it is feasible to increase the series resistance by x10 or not.
Some details:
Note that in this first computation, I have not considered
- Actuation range required by other loops (e.g. local damping, Oplev etc).
- At some point, I need to add the 2P/c radiation pressure disturbance as well.
- The control signal is calculated assuming we are actuating equally on both ETMs (but with opposite phase).
- RMS computation is done from 30 Hz downwards, as above 30 Hz, I think the estimate from the previous elog is not true seismic displacement.
- De-whitening filters (or digital whitening), which will be required to suppress DAC noise at 100Hz.
- DARM loop shape, specifically low-pass to avoid sensing noise injection. In this calculation, I just used the pendulum transfer function.
While doing this calculation, I have accounted for the fact that right now, the analog de-whitening filters in the ETM drive chain have a x3 gain which we will remove. Actually this is an assumption, I have not yet measured a transfer function, maybe I'll do one channel at EY to confirm. Also, the actuator gains themselves need to be confirmed.
As I was looking at the coil driver schematic more closely, I realized that there are actually two separate series resistances, one for the fast controls path, and another for the DC bias voltage from the slow ADCs. So I think we have been underestimating the Johnson noise of the coil drivers by sqrt(2). I've also attached screenshots of the IFOalign and MCalign screens. The two ITMs and ETMX have pitch DC bias values that are compatible with a x10 increase of the series resistance. But even so, we will have ~3pA/rtHz per coil from the two resistances.
gautam 8pm May8: Seems like I had confirmed the x3 gain in the EX de-whitening board when Johannes and I were investigating the AI board offset. |
Attachment 1: darmProj.pdf
|
|
Attachment 2: 37.png
|
|
Attachment 3: MCalign_20180507.png
|
|
13821
|
Mon May 7 15:27:28 2018 |
gautam | Update | SUS | Stack measurement expectation |
[steve,gautam]
We tried to estimate what the load cell measurement should yield. Here is the weight breakdown (fudge factor for Al table is to try and account for tapped holes):
Element
|
Diameter [m]
|
Height [m]
|
Density [kg/m^3]
|
Mass [kg]
|
Number or fudge factor
|
Dim in inches
|
Table |
1.22 |
0.08 |
2700.00 |
240.07 |
0.85 |
Dia=48", thickness=3" |
Stack leg |
0.36 |
0.13 |
8000.00 |
100.85 |
9 |
Dia=14", thickness=5" |
Base plate |
1.37 |
0.05 |
8000.00 |
600.18 |
1 |
Dia=60",thickness=2" |
Base rods |
0.10 |
1.83 |
8000.00 |
118.55 |
2 |
Dia=4", length=6ft |
Stuff on table |
|
|
|
100.00 |
|
|
Blue beams |
|
|
|
100.00 |
|
|
|
|
|
|
|
|
|
Total [kg] |
|
|
|
2149.01 |
|
|
Total [lbs] |
|
|
|
4835.28 |
|
|
- Steve pointed out that there is some material removed from the stack legs for stability (hollows into which the viton springs fit). These countersinks have dimensions of diameter=2", height=1.75". So if we assume each leg has 10% less mass, the total weight becomes ~4600lbs.
- I think we will need to use one more load cell (i.e. total 4) for this measurement (we have more load cells, just need to setup one more controller).
- Steve is looking into acquiring some low profile jacks to deal with the fact that we only have limited travel range on the overall stack height because of the bellows.
- A useful document, from which we pulled some numbers (which also look reasonable using estimated dimensions and density calculations): P952005
|
Attachment 1: 40m_TMstack.JPG
|
|
13820
|
Mon May 7 11:46:07 2018 |
gautam | Update | PEM | FW parameter update |
As part of investigation into this issue, Jonathan Hanks pointed out that the "minute trends" being recorded by our system were actually only being recorded every 120 seconds (a.k.a. 2 minutes). He had fixed the appropriate line in the parameter file, but had not restarted the FW processes. I had restarted it on Friday. (but failed to elog it !) 
To check if this made any difference, I pulled 1 hour of "minute trend" data for the PSL table temperature channel from ~1 hour ago, and compared the number of datapoints against a 1 hour minute trend time series from 2 May. I've put the display with # of datapoints (for an identical length of time) from before [left] and after [right] the restart next to the plots in Attachment #1. Seems like we are getting minute trends written every 60 seconds now, as it should be .
The unavailability of trends from nodus is a separate issue for which JH has suggested another fix, to be elogged separately.
Quote: |
for whatever reason, I am unable to get minute or second trends from nodus for any channels (IMC, PEM, etc) since the reboot. has there been some more recent FB failure or is this still a bug since last years FB catastrophe?
|
|
Attachment 1: FWreboot.png
|
|
13819
|
Sat May 5 22:32:07 2018 |
Koji | Update | PSL | Modulation depth measurement for the 3IFO aLIGO EOM |
The 3IFO EOM was formerly tuned as the H2 EOM, so the resonant frequencies are different from the nominal aLIGO ones.
PORT1: 8.628MHz / 101 +/- 6 mrad_pk/V_pk
PORT2: 24.082MHz / 41.2 +/- 0.7 mrad_pk/V_pk
PORT3: 43.332MHz / 62.2 +/- 4 mrad_pk/V_pk
9MHz modulation is about x2.4 better than the one installed at LHO.
24MHz modulation is about x14 better. (This is OK as the new 24MHz is not configured to be resonant.)
45MHz modulation is about x1.4 better.
|
13818
|
Sat May 5 20:30:21 2018 |
Koji | Update | PSL | Modulation depth measurement for the 3IFO aLIGO EOM and aftermath |
Caution: Because of this work and my negligence, the RF output of the main Marconi for the IFO modulation is probably off. The amplifier (freq gen. box) was turned on. Therefore, we need to turn the Marconi on for the IFO locking.
I worked on my EOM m easurement using the beat setup. As there was the aux injection electronics, I performed my measurement having tried not to disturb the aux setup. The aux Marconi, the splitted PD output, and an open channel of the oscilloscope were used for my purpose. I have brought the RF spectrum analyzer from the control room. I think I have restored all the electronics back as before. I have re-aligned the beat setup after the EOM removed. Note that the aux NPRO, which had been on, was turned off to save the remaining life of the laser diode. |
13817
|
Fri May 4 21:17:57 2018 |
gautam | Configuration | ALS | BeathMouth pulled out of PSL table |
I have been puzzled about the beat note level we get out of the BeatMouth for some time.
- The beat PD used is the Menlo FPD310.
- But the version we have is an obsolete version of the product, for which a manual is hard to find.
- Hence, I don't know the transimpedance/electrical characteristics of this PD.
- The optical damage threshold of the PD is quoted as 2mW, which is a number I have been careful not to exceed.
- But I've felt that the beat amplitude level we get out of this PD is far too low considering the amount of DC optical power (as measured with a fiber power meter) incident on the PD.
- After some emailing and online hunting, I've gathered some numbers for the PD which are now on the wiki.
- The fiber beam splitters we use inside the BeatMouth don't have PM fibers. There are 3 such splitters inside the BeatMouth. So the overlap efficiency on the PD is unknown.
- But even so, the beat levels I was seeing were too low.
I have pulled the box out in order to re-characterize the DC power levels incident on the PD, and also to change the gain setting on the PD to the lower gain which is more compatible with the level of optical power we have going into the BeatMouth. The modern catalog for the FPD310 (see wiki) suggests that the maximum output voltage swing of the PD is 1Vpp driving a 50ohm load. With 50% overlapping efficiency between the PSL and AUX beams, and 400 uW of optical power from each beam, I expect an output of 0.5Vpp. Even with perfect overlap, I expect 0.8Vpp. So these numbers seem reasonable.
I also plan to check the scaling of electrical beat amplitude to optical power for a couple of levels to see that these scale as expected... |
13816
|
Fri May 4 19:06:28 2018 |
rana | Configuration | Electronics | AUX-PSL PLL Implementation & Characterization |
this doesn't make much sense to me; the phase to frequency conversion (mixer-demod to PZT ) should give us a 1/f loop as Johannes mentioned in the meeting. That doesn't agree with your loop shape.
How about give us some more details of the setup including photos and signal/power levels? And maybe measure the LB1005 TF by itself to find out what's wrong with the loop. |
13815
|
Fri May 4 18:59:39 2018 |
gautam | Update | SUS | Stack measurement ongoing |
[SV,KA,RXA,GV]
The stack weight measurement is going on at EX. ETMX watchdog is shutdown. Area is off limits over the weekend until the test is finished.
Not related to this work, but the dog clamps used on the EX table have to be re-positioned such that the clamping force is better distributed. The 2" beam splitter mount used to pick off a portion of the EX NPRO beam to the fiber has to be rotated. Also, there was a M6.9 EQ in Hawaii while we were doing this work it seems.. |
13814
|
Fri May 4 13:24:56 2018 |
Jon Richardson | Configuration | Electronics | AUX-PSL PLL Implementation & Characterization |
Attached are final details of the phase-locked loop (PLL) implementation we'll use for slaving the AUX 700 mW NPRO laser to the PSL.
The first image is a schematic of the electronics used to create the analog loop. They are curently housed on an analyzer cart beside the PSL table. If this setup is made permanent, we will move them to a location inside the PSL table enclosure.
The second image is the measured transfer function of the closed loop. It achieves approximately 20 dB of noise suppression at low frequencies, with a UGF of 50 kHz. In this configuration, locks were observed to hold for 10s of minutes. |
Attachment 1: PLL_Schematic.pdf
|
|
Attachment 2: PLL_AUX-PSL_40m.pdf
|
|